OpenStudy (anonymous):

What is the 8th term of the geometric sequence 3, 9, 27, …?

OpenStudy (anonymous):

a1 = 3^1 a2 = 3^2 therefore a n = 3^n a 8 = 3^ 8 = 6561

OpenStudy (anonymous):

do you see the pattern?

OpenStudy (anonymous):

Thanks this is atually a pretest and I'm just learning how to do this and I was trying to find a simplier way then sitting there and just adding and subtracting forever